What is the frequency of a 700 nm wave through a diamond?

n = 2.4
c = 3 x10^8
Answer #1

Wavelength = 700×10^-9m

V = C/n = 3×10^8/2.4 = 1.25×10^8m/s

Frequency f = V/wavelength = 1.25×10^8/700×10^-9 = 0.1785×10^15Hz

=) f = 178.5×10^12Hz